OpenStudy (anonymous):

A line passes through the points (-10, -4) and (-1,2) what is the y- intercept

OpenStudy (solomonzelman):

first find the slope y1 - y2 -4 - 2 -6 ------ ---> ------ ---> ---- = 2/3 x1 - x2 -10 - -1 -9 so, you know so far y=(2/3)x+b the y-intercept is still unknown, plug any of the 2 points to find it. y=(2/3)x+b lets pick the first points, (-1,2) (2)=(2/3)(-1)+b 2=-2/3+b 2+2/3=b b=2, 2/3 So the y-intercept is 2, 2/3 or 8/3

OpenStudy (anonymous):

Solomon would it be (,0 8/3) 0r (8/3,0)

OpenStudy (solomonzelman):

I see what you are asking, y-intercept is at x=0 and intersects the y axis. y-intercept should look like ( 0 , y-intercept )